OpenStudy (anonymous):

1/8x>-3/16 what is the solution set

OpenStudy (anonymous):

We are given \[\frac{1}{8}x>-\frac{3}{16}\] Multiplying everything by 8, we receive: \[1x>-\frac{3(8)}{16}\] Which is \[x>-\frac{3}{2}\]Our solution set, is, then:\[x\in\left(-\frac{3}{2},\infty\right)\]

OpenStudy (anonymous):

thanks